# Semi-Supervised Preference Optimization with Limited Feedback

### v1

@@ 5.4 Latent Preference in the Unpaired Data: Case Study
Substring Matching. To identify the specific benefits of unpaired data in SSPO, we conducted a comprehensive analysis using a substring matching algorithm to find semantically similar content between AlpacaEval2.0 questions and our unpaired datasets. For each AlpacaEval2.0 question, we computed semantic similarity scores with both paired and unpaired data, identifying cases where questions had high similarity to unpaired content ($\geq 0.6$) but low similarity to paired data ($\leq 0.3$). This analysis revealed numerous instances where SSPO’s performance was enhanced by exposure to relevant unpaired data during training. Table 5 presents a concrete example where SSPO benefited from unpaired data exposure in the UltraChat. We show the AlpacaEval2.0 question, the most similar instruction and corresponding response in each unpaired data, and the answer to the question from fine-tuned models. SSPO successfully aligns its generation based on such high-quality unpaired responses, resulting in better evaluation outcomes. More case studies and qualitative examples are provided in Appendix D.1 and D.2, respectively.
@@ Appendix A Statement on the Use of Large Language Models
During the research, Large Language Models (LLMs) were utilized in a limited, supporting capacity. Their utilization was confined to post-drafting revisions, particularly for grammatical corrections, sentence refinement, and enhancing overall writing consistency. The core intellectual contributions, including the research ideas, methodology, data analysis, and conclusion are solely the work of the authors. No substantive or analytical content was generated by LLMs, ensuring that their application did not compromise the academic integrity and originality of this research.
@@ B.2 Gradient Analysis
We provide a detailed analysis for the gradient behavior of our objective, separately for the supervised (paired) and semi-supervised (unpaired) components. This analysis reveals how the model encourages correct ranking through reward-based learning signals, and how pseudo-labels influence parameter updates.
@@
We define our paired-data risk as Eq. (4). The gradient weight increases when the model incorrectly assigns a higher reward to the losing response. This encourages the model to correct such misrankings by increasing the winning score and decreasing the losing one. This analysis is basically the same as DPO (Rafailov et al., 2024) and SimPO (Meng et al., 2024).
@@
Here, $\beta$ is the same scaling coefficient used in the paired-data risk, and $P_{D_{U}}(s)$ is the prior for pseudo-labeling. The gradient direction encourages increasing the reward of confident winning samples ($\tilde{s}=1$) and decreasing the reward of confident losing samples ($\tilde{s}=0$), using the threshold $\hat{\delta}$ derived from the paired dataset.
This reflects how pseudo-labeled responses influence learning through contrastive updates with respect to a reward threshold. When a response is pseudo-labeled as winning (i.e., its reward exceeds the threshold), the model increases its reward by a factor proportional to the confidence $\sigma(\hat{\delta}-r_{\theta})$, encouraging the model to reinforce highly ranked unpaired responses. Conversely, when the response falls below the threshold and is labeled as losing, the gradient acts to decrease its reward, again modulated by a sigmoid-based confidence score. Notably, the threshold $\hat{\delta}$ serves as a stable anchor for all unpaired updates, effectively transferring alignment signals from supervised data to the unpaired setting. As a result, the unpaired objective promotes consistent reward separation even in the lack of explicit preference pairs.
@@ C.1 Implementation of Pseudo-Labeled Risk
Note that $\hat{p}_{w}^{B}$ and $\hat{p}_{l}^{B}$ares the estimated density of winning and losing rewards within a minibatch B, respectively. $\mu_{B}$ and $\sigma_{B}$ denote the mean and standard deviation of reward values for all responses within the minibatch. We set the momentum parameter $m=0.95$.
This formulation ensures that the pseudo-labeling threshold remains stable and robust throughout training, even when the reward distribution shifts dynamically. The combination of theoretically grounded threshold estimation via KDE and practical stabilization through EMA provides both theoretical soundness and training stability.
@@ C.2 Toy Experiment
We report full results of test accuracy onthe toy dataset without or with noise, with a fixed number of unpaired samples ($n_{U}$=1000) in Table 7. SSPO consistently and significantly outperforms all baselines across different quantities of labeled data ($n_{L}$). The advantage is most pronounced in the extremely data-scarce setting ($n_{L}$=10), a substantial improvement over the best baselines. This result underscores SSPO’s data efficiency, originating in its ability to effectively utilize the large pool of unlabeled data. Furthermore, SSPO’s performance is robust, consistently delivering notable results across a broad spectrum of prior probability settings for the unpaired data. Table 7 describes a detailed configuration of toy experiments. The training procedure follows standard practices with 10 epochs,a total batch sizeof 16, and model selection based on the minimum of validation loss.
@@ C.3 Real-Data Experiments
Table 8 describes a detailed configuration of experiments on real-world datasets. Our implementation builds upon the LlamaFactory framework Zheng et al. (2024), a widely used open-source library for training, fine-tuning, and serving LLMs. We detail the hyperparameter configurations used for each baseline method across the three backbone models. All experiments were conducted using 2 GPUs of A6000, with tuning strategies adjusted per model size. All unspecified hyperparameters for Mistral and Llama3 baselines follow the configurations provided in SimPO [^2] https://github.com/princeton-nlp/SimPO (Meng et al., 2024).
We use LoRA (Hu et al., 2021) for rank 8 after exploring ranks from the set {4, 8, 32, 256} for all backbones. We perform a learning rate sweep between 1e-4 and 1e-8 and report results with the best performing value. The total batch size, gradient accumulation steps, and cutoff length are fixed at 64, 8, and 1024, respectively, across all models. We select the best model based on the minimum validation loss.
For SSRM, we follow the original protocol by using a preset confidence threshold and generating response pairs from prompts with no annotation for pseudo-labeling after training the reward model [^3] https://github.com/RLHFlow/RLHF-Reward-Modeling. To ensure fair comparison, we generate one additional response using the SFT model of each backbone and use the reward model trained by SimPO at each data size. We run three iterations for each setting with a default confidence level of 0.8, meaning the model incorporates new data pairs only when they exceed this confidence threshold.
For SSPO, we fix $\beta=10$ and $\Delta=2$ and the adaptive scheduler $\gamma^{\prime}$ is bounded below by $\gamma_{\min}=n_{L}/(n_{L}+n_{U})$. We set the pseudo-labeling prior to 0.5 since there is no suitable preference information available for the UltraChat.
Table 8: Hyperparameter settings for real-data experiments across all domains and backbones. SSPO is configured with prior 0.5 under two different data sizes.
@@
Method Hyperparameter Phi-2 Mistral Llama3 Mistral Llama3 Mistral Llama3
DPO Training Epochs 1 1 1 1 1 1 1
@@
ORPO Training Epochs 1 1 1 1 1 1 1
@@
SimPO
SSRM
Training Epochs 1 1 1 1 1 1 1
@@
KTO Training Epochs 1 1 1 1 1 1 1
@@
SPA Training Epochs 1 1 1 1 1 1 1
@@
SSPO 1% data size
Training Epochs 2 2 2 1 1 3 3
@@
Decay Rate ($\lambda$) 0.005 0.001 0.01 0.01 0.001 0.001 0.001
10% data size
Training Epochs 1 1 2 1 1 1 1
@@ C.4 Pseudo-Code of SSPO
21:  $\mathcal{L}(f_{\theta})\leftarrow\gamma_{t}\cdot R_{B_{L}}(f_{\theta})+(1-m)\cdot R_{B_{U}}(f_{\theta})$
@@
Appendix D Qualitative Examples
D.1 More Case Studies

### v2

@@ 5.4 Latent Preference in the Unpaired Data: Case Study
Substring Matching. To identify the specific benefits of unpaired data in SSPO, we conducted a comprehensive analysis using a substring matching algorithm to find semantically similar content between AlpacaEval2.0 questions and our unpaired datasets. For each AlpacaEval2.0 question, we computed semantic similarity scores with both paired and unpaired data, identifying cases where questions had high similarity to unpaired content ($\geq 0.6$) but low similarity to paired data ($\leq 0.3$). This analysis revealed numerous instances where SSPO’s performance was enhanced by exposure to relevant unpaired data during training. Table 5 presents a concrete example where SSPO benefited from unpaired data exposure in the UltraChat. We show the AlpacaEval2.0 question, the most similar instruction and corresponding response in each unpaired data, and the answer to the question from fine-tuned models. SSPO successfully aligns its generation based on such high-quality unpaired responses, resulting in better evaluation outcomes. More case studies and qualitative examples are provided in Appendix G.1 and G.2, respectively.
@@ Appendix A Statement on the Use of Large Language Models
During the research, Large Language Models (LLMs) were utilized in a limited, supporting capacity. Their utilization was confined to post-drafting revisions, particularly for grammatical corrections, sentence refinement, and enhancing overall writing consistency. The core intellectual contributions, including the research ideas, methodology, data analysis, and conclusion, are solely the work of the authors. No substantive or analytical content was generated by LLMs, ensuring that their application did not compromise the academic integrity and originality of this research.
@@ B.2 Gradient Analysis
We provide a detailed analysis of the gradient behavior of our objective, separately for the supervised (paired) and semi-supervised (unpaired) components. This analysis reveals how the model encourages correct ranking through reward-based learning signals, and how pseudo-labels influence parameter updates.
@@
We define our paired-data risk as Eq. (4). The gradient weight increases when the model incorrectly assigns a higher reward to the losing response. This encourages the model to correct such misrankings by increasing the winning score and decreasing the losing one. This analysis is basically the same as DPO (rafailov2024direct) and SimPO (meng2024simposimplepreferenceoptimization).
@@
[…]
Combining the two components, the full gradient of our objective in Eq. (12) can be written as
$\nabla_{\theta}\mathcal{L}(f_{\theta})=\gamma^{\prime}\,\nabla_{\theta}R_{D_{L}}(f_{\theta})+(1-\gamma^{\prime})\,\nabla_{\theta}R_{D_{U}}(f_{\theta}).$
As discussed in Section 4.3, $\gamma^{\prime}$ starts from 11 and decays toward $\gamma_{\min}$, so that early updates are dominated by $\nabla_{\theta}R_{D_{L}}(f_{\theta})$ from human-labeled comparisons, while the influence of $\nabla_{\theta}R_{D_{U}}(f_{\theta})$ grows only after the reward model has already learned a meaningful separation between winning and losing responses. This interaction between the scheduler and the gradient structure explains why SSPO remains robust even when pseudo-labels are imperfect, as also evidenced by our toy experiments with heavy label noise and our domain-specific experiments in Section 5.2.
@@ C.1 Implementation of Pseudo-Labeled Risk
Note that $\hat{p}_{w}^{B}$ and $\hat{p}_{l}^{B}$ are the estimated density of winning and losing rewards within a minibatch B, respectively. $\mu_{B}$ and $\sigma_{B}$ denote the mean and standard deviation of reward values for all responses within the minibatch. We set the momentum parameter $m=0.95$.
This formulation ensures that the pseudo-labeling threshold remains stable and robust throughout training, even when the reward distribution shifts dynamically. From a theoretical perspective, the KDE-based update of $\hat{\delta}_{B}$ can be seen as a mini-batch approximation of the global Bayes-risk minimizer in Section 4.2, while the EMA statistics $(\mu_{t},\sigma_{t},\hat{\delta}_{t})$ smooth out stochastic fluctuations across batches. In the limit of many updates, this procedure tracks a slowly varying approximation to the population-level optimal threshold, while remaining numerically stable in the non-stationary regime induced by ongoing policy updates. The combination of theoretically grounded threshold estimation via KDE and practical stabilization through EMA provides both theoretical soundness and training stability.
@@ C.2 Toy Experiment
We report full results of test accuracy on the toy dataset without or with noise, with a fixed number of unpaired samples ($n_{U}$=1,000) in Table 7. SSPO consistently and significantly outperforms all baselines across different quantities of labeled data ($n_{L}$). The advantage is most pronounced in the extremely data-scarce setting ($n_{L}$=10), a substantial improvement over the best baselines. This result underscores SSPO’s data efficiency, originating in its ability to effectively utilize the large pool of unlabeled data. Furthermore, SSPO’s performance is robust, consistently delivering notable results across a broad spectrum of prior probability settings for the unpaired data. Table 7 describes a detailed configuration of toy experiments. The training procedure follows standard practices with 10 epochs, a total batch size of 16, and model selection based on the minimum of validation loss.
@@ C.3 Real-Data Experiments
Table 8 describes a detailed configuration of experiments on real-world datasets. Our implementation builds upon the LlamaFactory framework (zheng2024llamafactory), a widely used open-source library for training, fine-tuning, and serving LLMs. We detail the hyperparameter configurations used for each baseline method across the three backbone models. All experiments were conducted using 2 GPUs of A6000, with tuning strategies adjusted per model size. All unspecified hyperparameters for Mistral and Llama3 baselines follow the configurations provided in SimPO [^2] https://github.com/princeton-nlp/SimPO (meng2024simposimplepreferenceoptimization).
Table 8: Hyperparameter settings for real-data experiments across all domains and backbones. All the algorithms are configured with prior 0.5 under two different data sizes.
@@
Method Size Hyperparameter Phi-2 Mistral Llama3 Mistral Llama3 Mistral Llama3
DPO 1% Training Epochs 1 1 1 1 1 1 1
@@
10% Training Epochs 1 1 1 1 1 1 1
Learning Rate 1e-6 5e-7 1e-6 5e-7 5e-7 5e-6 1e-6
$\beta$ 0.1 0.1 0.1 0.1 0.1 0.1 0.1
ORPO 1% Training Epochs 1 1 1 1 1 1 1
@@
10% Training Epochs 1 1 1 1 1 1 1
Learning Rate 7e-6 5e-6 5e-6 3e-7 3e-7 1e-6 5e-6
$\lambda$ 0.25 0.2 0.2 0.2 0.2 0.25 0.2
SimPO 1% Training Epochs 1 1 1 1 1 1 1
@@
10% Training Epochs 1 1 1 1 1 1 1
Learning Rate 5e-7 5e-7 1.5e-7 5e-7 5e-7 5e-7 6e-7
$\beta$ 10 2 2 2 2 2 2
$\Delta$ 2 0.8 1 0.8 0.8 0.8 1
SSRM 1% Training Epochs 1 1 1 1 1 1 1
Learning Rate 5e-7 5e-7 6e-7 5e-7 5e-7 6e-7 6e-7
$\beta$ 10 2 2 2 2 2 2
$\Delta$ 2 0.8 1 0.8 0.8 1 1
10% Training Epochs 1 1 1 1 1 1 1
Learning Rate 5e-7 5e-7 6e-7 5e-7 5e-7 6e-7 6e-7
$\beta$ 10 2 2 2 2 2 2
$\Delta$ 2 0.8 1 0.8 0.8 1 1
KTO 1% Training Epochs 1 1 1 1 1 1 1
@@
10% Training Epochs 1 1 1 1 1 1 1
Learning Rate 5e-6 5e-6 5e-6 1e-6 1e-6 1e-6 1e-6
SPA 1% Training Epochs 1 1 1 1 1 1 1
@@
10% Training Epochs 1 1 1 1 1 1 1
Learning Rate 5e-6 2e-5 1e-5 1e-5 1e-5 2e-5 1e-5
$\beta$ 0.01 0.01 0.01 0.01 0.01 0.01 0.01
Iteration 1 2 2 2 2 2 2
SSPO 1% Training Epochs 2 2 1 1 1 3 3
@@
Decay Rate ($\lambda$) 0.005 0.001 0.005 0.01 0.001 0.001 0.001
10% Training Epochs 1 1 2 1 1 1 1
@@
We use LoRA (hu2021loralowrankadaptationlarge) for rank 8 after exploring ranks from the set {4, 8, 32, 256} for all backbones. We perform a learning rate sweep between 1e-4 and 1e-9 and report results with the best performing value. The total batch size, gradient accumulation steps, and cutoff length are fixed at 64, 8, and 1024, respectively, across all models. We select the best model based on the minimum validation loss.
For SSRM, we follow the original protocol by using a preset confidence threshold and generating response pairs from prompts with no annotation for pseudo-labeling after training the reward model [^3] https://github.com/RLHFlow/RLHF-Reward-Modeling. To ensure fair comparison, we generate one additional response using the SFT model of each backbone and use the reward model at each data size. We run three iterations for each setting with a default confidence level of 0.8, meaning the model incorporates new data pairs only when they exceed this confidence threshold. For SSPO, we fix $\beta=10$ and $\Delta=2$ and the adaptive scheduler $\gamma^{\prime}$ is bounded below by $\gamma_{\min}=n_{L}/(n_{L}+n_{U})$. We set the pseudo-labeling prior to 0.5 since there is no suitable preference information available for the UltraChat.
@@ C.4 Pseudo-Code of SSPO
21:  $\mathcal{L}(f_{\theta})\leftarrow\gamma_{t}\cdot R_{B_{L}}(f_{\theta})+(1-\gamma_{t})\cdot R_{B_{U}}(f_{\theta})$
@@
Appendix D The Difference between SSRM, SPA, and SSPO
In this section, we provide a detailed comparison between SSPO and two prominent semi-supervised alignment frameworks: Semi-Supervised Reward Modeling (SSRM(he-etal-2024-semi-supervised) and Spread Preference Annotation (SPA(kim2025spreadpreferenceannotationdirect). While all three methods aim to leverage unlabeled data to mitigate the scarcity of human preference labels, SSPO distinguishes itself through its theoretical grounding, data efficiency, and non-iterative training paradigm. We analyze these methods across four key dimensions: Data Composition, Algorithmic Formulation, Theoretical Foundation, and Computational Efficiency.
D.1 Data Composition & Annotation
Let $x$ be a prompt (i.e., an input of our model $\pi_{\theta}$ to be trained) and $y$ be a corresponding response (i.e., an output of our model $\pi_{\theta}$). All three algorithms leverage both labeled and unlabeled datasets, but the way to formulate and generate the datasets is quite different.
  • SSRM: Requires a small set of paired and labeled data $D_{L}=\{(x^{(i)},y_{w}^{(i)},y_{l}^{(i)})\}_{i=1}^{n_{L}}$ and a large set of unlabeled pairs $D_{U}=\{(x_{u}^{(i)},y_{u1}^{(i)},y_{u2}^{(i)})\}_{i=1}^{n_{U}}$. It relies on the model’s confidence to pseudo-label these pairs, essentially performing self-training on an existing paired structure.
  • SPA: Utilizes a small seed of paired data $D_{L}=\{(x^{(i)},y_{w}^{(i)},y_{l}^{(i)})\}_{i=1}^{n_{L}}$ and assumes that there exists a set of unseen prompts $X_{t}=\{x_{u}^{(t,i)}\}_{i=1}^{n_{U}^{(t)}}$ per $t$-th iteration. It explicitly generates multiple responses for the unseen prompts during training to form new unlabeled pairs $D_{U}^{(t)}=\{(x_{u}^{(t,i)},y_{u1}^{(t,i)},y_{u2}^{(t,i)})\}_{i=1}^{n_{U}^{(t)}}$ per $t$-th iteration, which are then self-labeled using the model’s implicit reward by DPO (rafailov2024direct).
  • SSPO (Ours): Requires a small set of paired data $D_{L}=\{(x^{(i)},y_{w}^{(i)},y_{l}^{(i)})\}_{i=1}^{n_{L}}$ and a large pool of unpaired and unlabeled data $D_{U}=\{(x_{u}^{(i)},y_{u}^{(i)})\}_{i=1}^{n_{U}}$. Unlike SSRM, it does not require pre-existing pairs in $D_{U}$, and unlike SPA, it does not require generating new responses during training. SSPO uniquely assigns “winning” or “losing” pseudo-labels to single responses based on a learned reward threshold $\delta$, allowing for the direct utilization of standard SFT datasets.
D.2 Algorithm & Formulation
  • SSRM & SPA: Both methods employ an iterative cycle: Training $\rightarrow$ Inference/Generation $\rightarrow$ Pseudo-labeling $\rightarrow$ Filtering $\rightarrow$ Re-training. SSRM treats reward modeling as a classification task, while SPA uses the DPO objective for self-refinement. This iterative process is inherently complex and sensitive to the quality of intermediate models.
  • SSPO (Ours): SSPO frames preference learning as a Bayes-optimal classification problem. It optimizes a joint objective combining a standard preference loss on $D_{L}$ and a threshold-based binary classification loss on $D_{U}$: $\mathcal{L}(f_{\theta})=\gamma^{\prime}\cdot R_{D_{L}}(f_{\theta})+(1-\gamma^{\prime})\cdot R_{D_{U}}(f_{\theta})$ (Eq. (12)). By dynamically estimating the optimal threshold $\hat{\delta}$ via Kernel Density Estimation (KDE) and using an adaptive scheduler $\gamma^{\prime}$, SSPO achieves robust alignment in a single training stage without iterative loops.
D.3 Theoretical Foundation
  • SSRM & SPA: Primarily rely on empirical validation of self-training principles (e.g., noisy label learning, curriculum learning) without specific theoretical guarantees for the optimality of their pseudo-labeling strategies in the context of preference distributions.
  • SSPO (Ours): Provides a rigorous theoretical foundation (Theorem 1). We prove the existence of an optimal reward threshold $\delta^{*}$ that separates winning and losing response distributions with high probability under sub-Gaussian assumptions. This theoretical guarantee underpins our pseudo-labeling strategy, ensuring it is statistically principled rather than heuristic.
D.4 Computational Efficiency
  • SSRM & SPA: Computationally expensive due to their iterative nature. They require multiple rounds of inference over the entire unlabeled dataset (and generation for SPA), followed by re-training.
  • SSPO (Ours): Highly efficient single-stage training. The additional cost of KDE and threshold estimation is negligible compared to the overhead of iterative generation and inference. SSPO reuses existing SFT datasets directly, eliminating the need for costly response generation during the alignment phase. We provide more analyses of time complexity in Appendix E.
Appendix E Computational Overhead
E.1 Time Complexity of SSPO
In this section, we give a simple big-O analysis of the computational overhead introduced by SSPO. Let $D_{L}$ denote the paired dataset and $D_{U}$ the unpaired dataset. At each training step $t$, Algorithm 1 samples a minibatch $B_{L}\subset D_{L}$ and $B_{U}\subset D_{U}$ and computes rewards for all responses in $B_{L}\cup B_{U}$. Let $N_{\text{tok}}$ be the total number of tokens in the current batch (including instructions and all responses), and let $d$ be the model dimension. The dominant cost per step comes from the forward and backward passes of the transformer model, which scale as $\mathcal{O}\big(N_{\text{tok}}\cdot d^{2}\big)$ for a single causal decoder, or $\mathcal{O}\big(N_{\text{tok}}\cdot H\cdot d^{2}\big)$ if we make the number of attention heads $H$ explicit. When a separate reference model is used, this cost is increased by a constant factor (roughly $\times 2$), but the asymptotic order remains unchanged.
SSPO adds a few extra components on top of this base cost:
  • Reward standardization and EMA updates. For a batch of $|B_{L}|+|B_{U}|$ rewards, computing the batch mean and variance, and updating the exponential moving averages $\mu_{t}$ and $\sigma_{t}$ costs $\mathcal{O}(|B_{L}|+|B_{U}|)$.
  • KDE-based Bayes-risk threshold. Let $M=|B_{L}|$ be the number of labeled pairs in the batch, and let $G$ be the number of grid points used to evaluate the kernel density estimators (we use $G=200$ in practice). The complexity of fitting Gaussian KDEs for the winning and losing rewards and searching for the Bayes-risk-minimizing threshold is $\mathcal{O}(M^{2}+MG)$, which we can conservatively write as $\mathcal{O}(M^{2})$ with respect to the batch size.
  • Pseudo-labeling of unpaired responses. For each unpaired response $(x_{u},y_{u})\in B_{U}$, SSPO compares the normalized reward to the current threshold and computes a logistic-style loss. This step scales linearly in the number of unpaired items, i.e., $\mathcal{O}(|B_{U}|)$.
Putting these terms together, the per-step complexity of SSPO can be summarized as
$\mathcal{O}\big(N_{\text{tok}}\cdot d^{2}\big)\;+\;\mathcal{O}(|B_{L}|+|B_{U}|)\;+\;\mathcal{O}(M^{2})\;\approx\;\mathcal{O}\big(N_{\text{tok}}\cdot d^{2}\big),$
where the transformer forward/backward passes clearly dominate. Over $\mathcal{T}$ training steps with a total of $N_{\text{tok}}$ processed tokens, the overall complexity becomes
$\mathcal{O}\big(N_{\text{tok}}\cdot d^{2}\big).$
In other words, SSPO has the same asymptotic time complexity as DPO and SimPO; the KDE-based thresholding and pseudo-labeling introduce only a modest constant-factor overhead that is negligible compared to the cost of running LLMs. Likewise, DPO incurs a slightly larger constant factor due to the additional reference-model forward pass, and both SimPO and SSPO yield the smallest constant factor because they require only a single policy-model forward/backward pass without any reference model.
E.2 Empirical Training Time of SSPO
We also report the empirical training time required to run SSPO under our experimental setup. Table 9 summarizes the observed training time per epoch and the average number of samples processed per second. All backbones are trained with LoRA adapters with rank 8 on 2 GPUs of A6000 with a total batch size of 64. Combined with the complexity analysis in Appendix E.1, these measurements indicate that SSPO is practically feasible.
Table 9: Per-epoch training runtime (Runtime) and training samples per second (Samples/s) of SSPO across backbones and paired-data sizes. UltraChat is always used at 10% ($n_{U}$=20,786) of its dataset, while UltraFeedback is used at either 1% ($n_{L}$=611) or 10% ($n_{L}$=6,113).
Size Phi-2 (2.7B) Mistral (7B) Llama3 (8B)
Runtime Samples/s Runtime Samples/s Runtime Samples/s
1% 4,764s (1.3h) 4.47 9,684s (2.7h) 2.20 9,282s (2.6h) 2.29
10% 7,813s (2.2h) 3.36 19,350s (5.4h) 1.35 15,919s (4.4h) 1.65
Appendix F More Quantitative Analyses
F.1 DPO and SimPO combined with Unpaired Data
For further comparison, we extend the preference optimization framework to two practical variants that jointly leverage paired and unpaired data. We combine a standard preference objective on paired data with a supervised fine-tuning (SFT) objective on unpaired data to both DPO and SimPO objectives to compare the performance with the same setting of data utilization. We call these variants DPO+SFT and SimPO+SFT. Both methods follow the same data formulation as SSPO but differ in the formulation of loss.
Setup. We denote by $\pi_{\theta}(y\mid x)$ the policy model and by $\pi_{\mathrm{ref}}(y\mid x)$ a fixed reference model when required. Given a pair $(x,y_{w},y_{l})$ of the paired dataset $D_{L}$, the DPO objective can be written as
$\mathcal{L}_{\mathrm{DPO}}(\theta;x,y_{w},y_{l})$ $=-\log\sigma\!\Big[\beta\big(\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\mathrm{ref}}(y_{w}\mid x)}-\log\frac{\pi_{\theta}(y_{l}\mid x)}{\pi_{\mathrm{ref}}(y_{l}\mid x)}\big)\Big],$
and the corresponding expected risk over $D_{L}$ is $R_{D_{L}}^{\mathrm{DPO}}(\theta)=\mathbb{E}_{(x,y_{w},y_{l})\sim D_{L}}\big[\mathcal{L}_{\mathrm{DPO}}(\theta;x,y_{w},y_{l})\big]$.
To incorporate abundant unpaired data $D_{U}$, we define a supervised fine-tuning loss that treats each $(x_{u},y_{u})$ as winning sample:
$\mathcal{L}_{\mathrm{SFT}}(\theta;x_{u},y_{u})=-\log\pi_{\theta}(y_{u}\mid x_{u}),$
and the corresponding unpaired risk is $R_{D_{U}}^{\mathrm{SFT}}(\theta)=\mathbb{E}_{(x_{u},y_{u})\sim D_{U}}\big[\mathcal{L}_{\mathrm{SFT}}(\theta;x_{u},y_{u})\big]$. The DPO+SFT objective is then defined as a weighted sum of the preference risk on $D_{L}$ and the SFT risk on $D_{U}$:
$\mathcal{L}_{\mathrm{DPO+SFT}}(\theta)$ $=R_{D_{L}}^{\mathrm{DPO}}(\theta)+\lambda_{\mathrm{SFT}}\cdot R_{D_{U}}^{\mathrm{SFT}}(\theta),$
where $\lambda_{\mathrm{SFT}}\geq 0$ controls the relative importance of the unpaired dataset term. When $\lambda_{\mathrm{SFT}}=0$, the objective reduces to standard DPO.
Similarly, using the reward function with length normalization and a prescribed margin $\Delta$, the SimPO loss on a single pair $(x,y_{w},y_{l})$ is
$\mathcal{L}_{\mathrm{SimPO}}(\theta;x,y_{w},y_{l})$ $=-\log\sigma\Bigg[\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}\mid x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}\mid x)-\Delta\Bigg],$
and the corresponding risk over $D_{L}$ is $R_{D_{L}}^{\mathrm{SimPO}}(\theta)=\mathbb{E}_{(x,y_{w},y_{l})\sim D_{L}}\big[\mathcal{L}_{\mathrm{SimPO}}(\theta;x,y_{w},y_{l})\big]$.
Analogously to DPO+SFT, we reuse the SFT loss on $D_{U}$ added with the length normalization:
$R_{D_{U}}^{\mathrm{SFT^{\prime}}}(\theta)=\mathbb{E}_{(x_{u},y_{u})\sim D_{U}}\bigg[-\frac{1}{|y_{u}|}\log\pi_{\theta}(y_{u}\mid x_{u})\bigg].$
Therefore, the SimPO+SFT objective is then given by
$\mathcal{L}_{\mathrm{SimPO+SFT}}(\theta)$ $=R_{D_{L}}^{\mathrm{SimPO}}(\theta)+\lambda_{\mathrm{SFT}}\cdot R_{D_{U}}^{\mathrm{SFT^{\prime}}}(\theta),$
using the same SFT coefficient $\lambda_{\mathrm{SFT}}$. This formulation yields a reference-free variant that still benefits from unpaired data.
Table 10: Comparing the performance of DPO+SFT and SimPO+SFT with SSPO at UltraFeedback. We set $\lambda_{\mathrm{SFT}}=1$ for both DPO+SFT and SimPO+SFT. The best numbers are in bold, and the second-best ones are underlined.
Baseline Size Mistral (7B)
LC WR MT
DPO+SFT 1% 26.5 17.0 7.6
10% 29.5 19.5 7.7
SimPO+SFT 1% 25.1 16.6 7.6
10% 27.3 17.3 7.6
SSPO 1% 26.7 18.1 7.7
10% 30.0 20.7 7.7
Results. Table 10 presents the comparative results on the UltraFeedback dataset. While the SFT-augmented baselines show improvements over their standard counterparts by leveraging the additional unpaired data, SSPO consistently outperforms all variants across different data sizes. This performance gap highlights the critical advantage of our algorithmic design. DPO+SFT and SimPO+SFT treat the unpaired data $D_{U}$ merely as a collection of positive demonstrations, applying a uniform SFT objective to all samples. This approach essentially regularizes the policy towards the distribution of the unpaired corpus but fails to distinguish the varying quality or preference alignment within that data. In contrast, SSPO does not indiscriminately reinforce all unpaired responses. By employing a dynamically learned reward threshold, SSPO actively performs pseudo-labeling, effectively filtering and differentiating pseudo-winning responses from pseudo-losing ones. This mechanism allows the model to learn a more nuanced preference boundary rather than simply imitating the aggregate statistics of the unpaired dataset. Therefore, the superior performance of SSPO is not merely a result of data scale, but a direct consequence of its ability to extract and leverage latent preference signals through principled semi-supervised learning.
F.2 Performances with Much Paired Data
We further investigated the performance of SSPO in regimes where paired preference data is abundant, to determine if the benefits of semi-supervised learning persist when the scarcity constraint is lifted. We conducted experiments on both the synthetic toy dataset and the real-world UltraFeedback dataset by scaling up the amount of paired data.
Table 11: Comparison of test accuracy on toy dataset without noise in different paired data. The best numbers for each size of $n_{L}$ are in bold, and the second-best ones are underlined.
Method $n_{L}=$ 10 $n_{L}=$ 50 $n_{L}=$ 100 $n_{L}=$ 1,000 $n_{L}=$ 10,000
DPO 0.743 0.777 0.846 0.986 0.995
ORPO 0.590 0.679 0.710 0.832 0.934
SimPO 0.762 0.776 0.817 0.981 0.996
SSPO (0.1) 0.822 0.910 0.965 0.986 0.992
SSPO (0.3) 0.810 0.889 0.955 0.983 0.993
SSPO (0.5) 0.841 0.879 0.960 0.983 0.996
SSPO (0.7) 0.800 0.857 0.966 0.985 0.991
SSPO (0.9) 0.840 0.861 0.965 0.986 0.991
Table 11 shows the test accuracy on the toy dataset as the number of labeled pairs ($n_{L}$) increases to 1.000 and 10,000. As expected, all methods converge towards high accuracy as the volume of supervision increases. This confirms that the $R_{D_{U}}(f_{\theta})$ does not conflict with $R_{D_{L}}(f_{\theta})$ but rather complements it, ensuring stability and high performance regardless of the size of paired data.
Table 12: Comparing the performance when training with 100% of UltraFeedback. The best numbers are in bold, and the second-best ones are underlined.
Baseline Size Mistral (7B)
LC WR MT
SPA 1% 18.2 15.6 7.7
10% 19.1 18.7 7.8
100% 26.2 19.3 7.7
SSPO 1% 26.7 18.1 7.7
10% 30.0 20.7 7.7
100% 32.4 21.0 7.8
In addition, Table 12 presents the results on the full UltraFeedback dataset (100% of data, $n_{L}$=61,135). Notably, SSPO trained on the complete labeled dataset achieves LC of 32.4%, significantly outperforming the strong semi-supervised baseline SPA (26.2%) trained on the same amount of labeled data. Furthermore, this performance is superior to SSPO trained on 10% data (30.0%), indicating that the model continues to benefit from the additional labeled data. This result suggests that even when ground-truth labels are sufficient, the principled pseudo-labeling of unpaired data provides signals that further refine the policy beyond what is possible with supervised signals alone.
Appendix G Qualitative Examples
G.1 More Case Studies